OpenStudy (anonymous):

we need 1 plank to redice the velocity of a bullet by 1/20 th part of it. now how many planks do we need to stop a bullet..?

OpenStudy (anonymous):

Of course, the usual formula relating wavelength (λ) and frequency (f = 1/T) for waves applies to transverse waves: v = s T μ = λ/T = λf (6) 2